### v1

# Inhibition of bacterial growth by antibiotics

Abstract
Growth in bacterial populations generally depends on the environment (availability and quality of nutrients, presence of a toxic inhibitor, product inhibition..). Here, we build a general model to describe the action of a bacteriostatic antibiotic, assuming that this drug inhibits essential autocatalytic cycles involved in the cell metabolism. The model can describe various types of antibiotics and confirms the existence of two distinct regimes of growth-dependent susceptibility, previously identified only for ribosome targeting antibiotics. Interestingly, below a certain threshold, a coexistence of two values of the growth rate is possible, which has also been observed experimentally.
autocatalytic networks, environment, antibiotics, growth
Introduction
The emergence of antibiotic resistance, which often occurs under changing levels of antibiotics is a major concern for human health [1]. In an important class of antibiotics, called bacteriostatic antibiotics [2], the drug does not induce death directly, but only renders some essential process in the cell metabolism less efficient or inactive [3, 4, 5, 6, 7, 8]. For these antibiotics, it thus appears essential to properly model cell metabolism and cell growth in order to better understand the action of antibiotics [9, 10, 11, 12].
In the field of bacterial growth, the experimental discovery of growth laws in the last decade [13, 12, 14, 15] represents a major step forward in our understanding of cell growth. These growth laws result from mass conservation and flux balance at steady-state. The first growth law has been derived using a comprehensive model of the cell metabolism based on the coupling of essential autocatalytic cycles, such as the cycle of ribosome production and that of RNA polymerase production [16]. This approach has also been used recently to formulate predictions about the interplay between cellular growth rate and mRNA abundances [17].
While predictions about the action of RNA-polymerase targeting antibiotics have also been derived from this framework, the full consequences for the inhibition of growth by a general antibiotics have not. In particular, Ref. [16] does not discuss the second growth law, nor the two modes of action of antibiotics, called reversible and irreversible binding regimes of antibiotics. This distinction is quite important in practice because for reversible binding, faster growth in the absence of the drug leads to an increased susceptibility, while the opposite is true for irreversible binding [12]. Further, the coexistence of two values of growth rate (growth rate bistability [18]) may occur below a certain threshold. At the moment, it is not known whether these behaviors should be expected for all types of antibiotics.
To summarize, we believe that the inhibition of bacterial growth by antibiotics has not been considered from a sufficiently general point of view, which is the approach we develop in the present paper. By building on Ref. [16], we develop a framework to describe the inhibition of bacterial growth by bacteriostatic antibiotics based on a model of cell metabolism. We first present our model, explore some of its consequences and extensions, and then show that it can describe successfully the dependence of the growth rate as function of the concentration of antibiotics for a wide range of different antibiotics. Further, we show that in a some limit, our general autocatalytic model allows to recover the equations of [12].
Model
We model the cell metabolism as two coupled autocatalytic cycles, in which one cycle describes the production of ribosomes, while the other describes RNA-polymerase production. These two autocatalytic cycles are coupled because ribosomes are necessary to synthesize RNA-polymerase protein subunits and vice-versa for ribosomes. To that basic model, we then add interaction with bacteriostatic antibiotics, as shown in the chemical network of Fig.1a: $B_{1}$ represents the number of active ribosomes; $C_{1}$ the number of active RNA polymerases; similarly $B_{2},...,B_{N-1}$ and $C_{2},...,C_{K-1}$ are the abundances of intermediates involved in the assembly of ribosomes and RNA polymerases respectively, $B_{N}$; $C_{K}$ are the abundances of fully assembled but resting ribosomes/RNA polymerases respectively, $R_{N},R_{K}$ are the abundances of building blocks needed to build $B_{N}$ and $C_{K}$.
Figure 1: (a) Scheme of coupled autocatalytic networks interacting with a toxic agent. The straight line linking two arrows represents a coupling through a min function [19, 20]. (b) Illustration of the growth laws when varying either the amount of antibiotics or the nutrient quality linked to pre-exposure growth rate $\lambda_{0}$ displayed on the right scale.
We suppose that "toxic" inhibiting agents in numbers $A$ can bind to one of the autocatalysts (chosen here to be $B_{1}$ for simplicity) with a rate $k_{on}$ and unbind with a rate $k_{off}$, proportionally to the relative abundance of antibiotics in the cell [12, 21]. We denote $B_{1,u}$ the abundance of unbound ribosomes and $B_{1,b}$ the abundance of bound ribosomes. The binding only occurs inside the cell, viewed as a compartment, in which antibiotics enter with rate $P_{in}$ and exit with rate $P_{out}$ (thanks to diffusion by passive transport or through pores by active transport) [22, 23]). The concentration of antibiotics outside the cell is $a_{ex}$. We use fractions measured with respect to the total number of mature individuals $B_{tot}=B_{1,u}+B_{1,b}+B_{N}$, assuming the total density of ribosomes remains constant in the cell of volume [24].
We rely on Leontief’s approach [19], or Liebig’s model in ecology [20], in which the rates of reactions involving two complementary resources are set by the limiting quantity among the two using a minimum function [16]. We denote $\tau_{life}$ the life time of mature individuals $B_{N},B_{1,u},B_{1,b}$, and we assume that we can neglect the inverse lifetimes of the intermediates $B_{2},...,B_{N-1}$. The names of the rates are self-explanatory and correspond to the transitions displayed in Fig.1a. In the following, we assume the cycle targeted by the toxic agent becomes limiting. Consequently, we can isolate this cycle and study its growth, because it restricts the growth of the rest of the network; the influence of the inhibition of the first cycle on the second cycle is studied in the Supplementary Material [21].
Due to balance growth of the cell, all species grow at the same rate $\lambda=d\ln{{\mathcal{N}}}/dt$, where $\mathcal{N}$ is typically the number of ribosomes or RNA-polymerases… One can then combine the equations of the model to obtain a linear matrix equation for the sub-populations of ribosomes only, without explicit dependence on antibiotics, and a self consistent equation for the growth rate $\lambda$ of the whole cycle (see Supplementary Material [21]). In the absence of inhibitors, the pre-exposure or basal growth rate is $\lambda_{0}$, which corresponds to the normal behaviour of the cell. As the concentration of antibiotics increases, the growth rate always decreases below this basal growth rate.
A key quantity is the fraction $Q(\lambda)=B_{1,u}/B_{tot}$, which takes the form of a polynomial :
$Q(\lambda)=\frac{1}{k_{B,1}}\left(1+\frac{\lambda}{k_{B,2}}\right)\times\ldots\times\left(1+\frac{\lambda}{k_{B,N-1}}\right)\left(\lambda+\frac{1}{\tau_{life}}\right).$ (1)
This polynomial simplifies in the limit of "fast assembly", which corresponds to $k_{B,2},...,k_{B,N-1}\gg\lambda$. In this case, we find the linear behaviour
$\frac{B_{1,u}}{B_{tot}}\simeq\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right).$ (2)
This linear increase between the fraction of unbound ribosomes and $\lambda$ is the first growth law [15, 13, 12, 16], which has been recovered from an autocatalytic cycles description of the cell in [16]. The law implies that by increasing the nutrient quality, which amounts to increase $\lambda_{0}$ by acting on assembly rates at external antibiotics concentration $a_{ex}=0$ with all other parameters fixed, we increase the growth rate and the fraction of mature ribosomes as illustrated in Fig. 1b.
The second growth law [13] predicts an increase of the growth rate together with a decrease in the activated ribosome fraction $(B_{1,u}+B_{1,b})/B_{tot}$ when the level of inhibition is increased. With our formalism, we obtain (see Supplementary Material [21]), in the limit of fast assembly, fast activation, long lifetime with a single intermediate step ($N=3$) :
$\frac{B_{1,u}+B_{1,b}}{B_{tot}}=1-\frac{\lambda}{k_{B,3}}.$ (3)
As shown on Fig.1b, by increasing the external concentration of antibiotics $a_{ex}$ with all other parameters fixed, we indeed predict an increase in the fraction of mature ribosomes and a decrease of the growth rate.
It is important to appreciate that the first and the second growth laws are derived from the model, they are not introduced as phenomenological constraints as done in Ref. [12]. It is also interesting to note that the solid blue curve in Fig. 1, which represents the limiting value of the ribosome fraction when $a_{ex}=0$ is not a linear function of $\lambda$, but a non-linear function with a negative curvature. This curvature was not expected in the original work on the growth laws [15] but it is also present in more detailed models of cell metabolism that include stochastic effects [25, 26].
We now explore further consequences of our formalism. For ribosomes, we can expect a long lifetime $\lambda\gg\frac{1}{\tau_{life}}$, a small resting rate $\lambda\gg k_{B,4}$, fast assembly and fast activation $\lambda\ll k_{B,3}$ [16]. These conditions translate to $\frac{1}{\tau_{life}},k_{B4}\ll\lambda_{0},k_{B1}\ll k_{B2},...,k_{B,N}$, yielding $\lambda_{0}\simeq k_{B1}$. In this limit, we can simplify our self-consistent equation for the growth rate :
$\frac{P_{in}a_{ex}}{\left(\frac{k_{B1}}{k_{on}}\frac{\lambda+P_{out}}{\lambda}+\frac{\lambda}{\lambda+k_{off}}\right)}\simeq\left(1-\frac{\lambda}{\lambda_{0}}\right)\left(\lambda+k_{off}\right).$ (4)
This equation is similar to that found in [12], which sets the growth rate of a bacteria in the presence of a bacteriostatic antibiotic. With the additional assumption of fast binding $\lambda_{0}\ll k_{on}$, the possible values of the growth rate are roots of a polynomial, from which it is possible to recover the reversible and irreversible limits of antibiotics binding [12]. In particular we find (see Supplementary Material [21]) a condition defining the transition from irreversible to reversible.
Reversible limit
The reversible limit corresponds to a strong outflux of toxic agents and unbinding rate. We find an equation setting the growth rate in this limit (see Supplementary Material [21]):
$Q(\lambda)=\frac{1}{1+\frac{K_{D}P_{in}}{P_{out}}a_{ex}}.$ (5)
In the limit $P_{out},k_{off}\gg\lambda_{0}$, the expression of the growth rate is that of Ref. [12], $\lambda=\frac{\lambda_{0}}{1+\frac{\lambda_{0}P_{in}a_{ex}}{k_{B,1}K_{D}P_{out}}}$, where $K_{D}$ is the dissociation constant $k_{off}/k_{on}$.
Irreversible limit
On the contrary, the irreversible limit corresponds to negligible outflux and unbinding rate compared to the influx of toxic agents and binding rate. Then, we obtain a different equation setting the growth rate (see Supplementary Material [21]):
$Q(\lambda)=1+\frac{P_{in}a_{ex}}{\lambda}.$ (6)
This equation typically has several solutions depending on the order of the polynomial $Q(\lambda)$. In the limit $P_{out},k_{off}\ll k_{on},\lambda_{0}$, we recover $\lambda=\frac{\lambda_{0}}{2}\left(1+\sqrt{1-\frac{4P_{in}a_{ex}}{\lambda_{0}}}\right)$ [12].
Interestingly, the self-consistent equation for the growth rate obtained within the autocatalytic framework (see Supplementary Material [21]) has two solutions in the irreversible limit with fast assembly, leading to two separate branches of solutions for $\lambda$. A first solution remains close to 00, corresponding to a non-growing cell. A second one is larger but exists only until a given concentration of inhibitors is reached, above which the system jumps on the other branch, and the growth rate vanishes as shown in Fig. 3a. In experiments, in the irreversible case, the system usually starts from $\lambda_{0}$ and the growth rate decreases as the concentration of inhibitors increases, until the discontinuity where the growth rate jumps on the second branch and vanishes. This growth bistability happens above a threshold, which can be determined by an implicit equation [21]. Such a phenomenon has been predicted in other theoretical work [12, 27], and it has also been observed experimentally [18, 26].
We have tested our model on a number of bacteriostatic antibiotics [2, 3, 6]: Chloramphenicol inhibits ribosome production by binding to ribosomes, preventing them from transcribing new proteins; Rifampicin targets RNA-polymerase by binding to RNA-polymerase [28, 29]; Kanamycin, Streptomycin, Chloramphenicol and Erythromycin target the ribosomal autocatalytic cycle [30, 3, 5, 7]; and finally Triclosan targets the synthesis of fatty acids [31, 32, 33], thus affecting the building of bacterial membranes [16].
Figure 2: Comparison with experiments for two bacteriostatic drugs, namely (a) Chloramphenicol (data from [8]) and (b) Kanamycin (data from [12]). The solid line shows the growth rate as a function of the fraction of inhibitors, while the dotted line shows a measure of the risk faced by the cell defined in the text. The data were fitted by constraining the parameters as explained is Supplementary Material [21].
In Fig.2, we show the normalized growth rate $\lambda/\lambda_{0}$ as function of the concentration of antibiotics only for Chloramphenicol and Kanamycin, the plots for the other antibiotics are shown in Supplementary Material [21].
In [16], the effects of Triclosan and Rifampicin were explained by adding Hill functions heuristically to describe saturation effects in the cycle. In contrast here, we provide an explicit expression for the dependence of the growth rate on the fraction of bacteriostatic antibiotics without such an assumption. The fact that we are able to describe a large panel of bacteriostatic antibiotics suggests that these antibiotics can indeed be depicted as inhibitors affecting essential cellular autocatalytic cycles despite their different mechanisms. Note that we recover different concavities in Fig.2, which correspond to the two distinct regimes of cellular response to the antibiotics previously identified for ribosome-targeting antibiotics [12]: the reversible limit where the outflux of antibiotics compensates the influx of the latter, and the irreversible limit where antibiotics bind quickly to autocatalysts, resulting in an accumulation of bound, inhibited individuals.
All these antibiotics are bacteriostatic agents, which slow growth but do not to induce death directly [11]. However, if the inhibition is too strong, processes that are necessary for survival cannot be satisfied and cell death can be induced in this way [34, 9]. To quantify this, we have introduced a measure of the risk faced by the cell, which we define as the fraction of bound active individuals $B_{1,b}$ with respect to unbound active individuals $B_{1,u}$ (see [21] for more details). The main interest of this notion is that it is independent of the type of action of the antibiotic and can be used to compare the efficiency of different antibiotics. This risk shown in Fig.2 as dashed lines is an increasing function of the concentration of antibiotics.
Half-inhibition concentration
The half-inhibition concentration $IC_{50}$ is defined as the concentration of toxic agent at which the growth rate is half its initial value. This is a measure of the sensitivity of the system to external stress, the higher it is, the more resistant is the system to inhibitors. We obtain an explicit expression for $IC_{50}$ in the limit of long lifetime and fast assembly, when the network contains an arbitrary number of steps $N$ (see [21] for details). If we can lump all intermediates into just one ($N=3$), we obtain
$\frac{IC_{50}}{IC_{50}^{*}}=\frac{1}{2}\left(\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+2K_{D}\frac{\lambda_{0}}{\lambda_{0}^{*}}\right)\left(1+\frac{\lambda_{0}}{2k_{off}}\right)+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right),$ (7)
where we have rescaled the half-inhibition concentration by a typical concentration $IC_{50}^{*}$ and the basal growth rate by a typical value $\lambda_{0}^{*}$ [21]. Note that this expression does not depend only on the ratio $\lambda_{0}/\lambda_{0}^{*}$ but also on $\lambda_{0}$ (itself defined by the parameters of the system). The rescaled half-inhibition concentration as a function of the rescaled basal growth rate in this limit is the convex function shown in Fig.3b. Remarkably, this function allows to collapse the measurements of many types of antibiotics. We reproduce in this figure experimental data from Ref. [12].
Additionally, we find in the limit of long lifetime, fast binding, fast assembly, and with $k_{off}\gg\lambda_{0}$, the rescaled half-inhibition concentration is essentially $\frac{IC_{50}}{IC_{50}^{*}}\simeq\frac{1}{2}\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right)$.
Figure 3: (a) Normalized growth rate versus the normalized antibiotic concentration. In dotted lines we represent the reversible regime $k_{off},P_{out}\geq k_{on},P_{in}$, in full lines the irreversible regime $k_{off},P_{out}\ll k_{on},P_{in}$. For the irreversible case (full lines), we observe two branches that represent the coexistence of two values of the growth rate, a "large" growth rate and a "near-zero" growth rate. A discontinuity appears when the system jumps from one branch to another. The colors of the curves correspond to different choices of rate constant $k_{B1}$ as shown on the scale on the right. $k_{B,1}$ essentially sets the basal growth rate $\lambda_{0}$ [21] and may vary from one cell to another in a population [35].
(b) Half-inhibition concentration $IC_{50}$ as function of the normalized pre-exposure growth rate in the case of no intermediate steps $m=0$. Symbols represent experimental data points extracted from Ref. [12], which correspond to various antibiotics as shown in the legend.
We recover in Fig.3b the two regimes of antibiotics binding mentioned above, the reversible regime where the half-inhibitory concentration decreases with $\lambda_{0}$ and the irreversible regime where it increases with $\lambda_{0}$. Adding intermediate steps shifts the minimum of the parabola towards lower $\lambda_{0}$ and reduces $IC_{50}$ and thus makes it easier to inhibit growth in the cycle. It also introduces a strong dependence of $IC_{50}$ on the rate constants $k_{1,B}$ in the reversible regime. This reflects that intermediate steps have a stronger impact in reversible pathways as compared to irreversible ones.
Conclusion
In this paper, building on previous works on cellular autocatalytic growth [16], we propose a general model for the inhibition of bacterial growth by antibiotics. This approach goes beyond Ref. [12] because growth laws are no longer introduced as additional constraints and an arbitrary number of steps is introduced in autocatalytic cycles. As we have shown, our model describes well the effects of a large panel of bacteriostatic antibiotics targeting key autocatalytic cycles in E.Coli. We have also found that the two regimes previously identified for ribosome-targeting antibiotics in [12], namely the reversible (strong outflux of inhibitors) and irreversible (small outflux of inhibitors) regimes, should in fact be expected generically for any bacteriostatic inhibitors targeting an autocatalytic cycle.
In the future, we would like to expand our approach towards bacteriocidal antibiotics, which are typically used in conjunction with bacteriostatic antibiotics in a time-dependent manner [36]. To understand cell death, one possibility would be to relate the measure of risk which we have introduced to the extinction probability of the cell. Experiments show significant cell-to-cell heterogeneity in antibiotic susceptibility [37], which require a model for the stochastic growth and death of individual cells and for the fluctuations in population size. In this respect, it is encouraging to see that our model predicts growth bistability, which could cause cell-to-cell heterogeneity, but clearly more work is needed to relate the single-cell and population susceptibility.
Finally, let us also point out that our approach based on autocatalytic cycles is rather general and could be applied beyond cellular biology to other fields, such as ecology [38] or economy, where individuals rather than molecules are able to create more of themselves thanks to autocatalytic cycles but can also be inhibited by toxic agents, either present in their environment or created by themselves as a result of their own growth.
Acknowledgements
We acknowledge inspiring discussions with C. Baroud and E. Maikranz, and insightful comments by L. Dinis and R. Pugatch.

### v2

@@
 We acknowledge inspiring discussions with C. Baroud and E. Maikranz, and insightful comments by L. Dinis and R. Pugatch.
+Appendix A Derivation of the growth laws in the simple case
+The chemical network we consider is shown on Fig. 1a of the main text. The role of the different species in the network is summarized in the following table
+$B_{1}$ Active ribosomes/Mature and active individuals
+$B_{2}$ Building blocks for ribosomes/Immature individuals
+$B_{3}$ Inactive ribosomes/ Mature and inactive individuals
+$C_{1}$ Active RNA-pol/Mature and active individuals
+$C_{2}$ Building blocks for RNA-pol/Immature individuals
+$C_{3}$ Inactive RNA-pol/ Mature and inactive individuals
+Table 1: Significance of the different species of the model
+According to Leontief’s approach [19], or Liebig’s model in ecology [20], the rates of reactions involving two complementary resources are set by the limiting quantity among the two using a minimum function as shown in Fig.
+1a.
+For this network, we can write the following equations :
+$\begin{split}\frac{dB_{1,u}}{dt}&=k_{B3}B_{3}-k_{B4}B_{1,u}-k_{on}\frac{A}{B_{tot}}B_{1,u}+k_{off}B_{1,b}-\frac{B_{1,u}}{\tau_{life}}\\ \frac{dB_{1,b}}{dt}&=k_{on}\frac{A}{B_{tot}}B_{1,u}-k_{off}B_{1,b}-\frac{B_{1,b}}{\tau_{life}}\\ \frac{dB_{2}}{dt}&=\min(k_{B1}B_{1,u},k_{C1}C_{1})-k_{B2}\min(B_{2},R_{1})\\ \frac{dB_{3}}{dt}&=k_{B2}\min(R_{1},B_{2})-k_{B3}B_{3}+k_{B4}B_{1,u}-\frac{B_{3}}{\tau_{life}}\\ \frac{dC_{1}}{dt}&=k_{C3}C_{3}-k_{C4}C_{1}\\ \frac{dC_{2}}{dt}&=\min(k_{B1}B_{1,u},k_{C1}C_{1})-k_{C2}\min(C_{2},R_{2})\\ \frac{dC_{3}}{dt}&=k_{C2}\min(R_{2},C_{2})-k_{C3}C_{3}+k_{C4}C_{1}\\ \frac{dR_{1}}{dt}&=P_{in,R_{1}}R_{1,ex}-k_{B2}\min(B_{2},R_{1})\\ \frac{dR_{2}}{dt}&=P_{in,R_{2}}R_{2,ex}-k_{C2}\min(C_{2},R_{2})\\ \frac{dA}{dt}&=P_{in}a_{ex}\Omega-P_{out}A-k_{on}\frac{A}{\Omega}B_{1,u}+k_{off}B_{1,b},\end{split}$ (8)
+where $\Omega$ is the volume of the cell, and $a_{ex}$ the external concentration of antibiotics.
+We assume that the total density of ribosomes $\rho$ remains constant [24], such that that $\Omega=B_{tot}/\rho$.
+In order to simplify notations, we absorb $\rho$ inside $P_{in}$ and $k_{on}$.
+When the species $B$ is limiting, we get a simpler system:
+$\begin{split}\frac{dB_{1,u}}{dt}&=k_{B3}B_{3}-k_{B4}B_{1,u}-k_{on}\frac{A}{B_{tot}}B_{1,u}+k_{off}B_{1,b}-\frac{B_{1,u}}{\tau_{life}}\\ \frac{dB_{1,b}}{dt}&=k_{on}\frac{A}{B_{tot}}B_{1,u}-k_{off}B_{1,b}-\frac{B_{1,b}}{\tau_{life}}\\ \frac{dB_{2}}{dt}&=k_{B1}B_{1,u}-k_{B2}B_{2}\\ \frac{dB_{3}}{dt}&=k_{B2}B_{2}-k_{B3}B_{3}+k_{B4}B_{1,u}-\frac{B_{3}}{\tau_{life}}\\ \frac{dA}{dt}&=P_{in}B_{tot}a_{ex}-P_{out}A-k_{on}\frac{A}{B_{tot}}B_{1,u}+k_{off}B_{1,b}.\end{split}$ (9)
+Now taking exponential solutions with growth rate $\lambda$, we obtain:
+$\begin{split}\left(\lambda+\frac{1}{\tau_{life}}+k_{on}\frac{A}{B_{tot}}+k_{B4}\right)B_{1,u}&=k_{B3}B_{3}+k_{off}B_{1,b}\\ \left(\lambda+\frac{1}{\tau_{life}}+k_{off}\right)B_{1,b}&=k_{on}AB_{1,u}\\ \left(\lambda+k_{B2}\right)B_{2}&=k_{B1}B_{1,u}\\ \left(\lambda+\frac{1}{\tau_{life}}+k_{B3}\right)B_{3}&=k_{B2}B_{2}+k_{B4}B_{1,u}\\ \left(\lambda+P_{out}+k_{on}\frac{B_{1,u}}{B_{tot}}\right)A&=P_{in}B_{tot}a_{ex}+k_{off}B_{1,b}.\end{split}$ (10)
+If we express everything in terms of the ratio to the total quantity of mature $B$, $B_{tot}=B_{1,u}+B_{1,b}+B_{3}$, $a=\frac{A}{B_{tot}}$, $b_{1,2,3}=\frac{B_{1,2,3}}{B_{tot}}$, we find by summing equations 1, 2 and 4:
+$\begin{split}\left(\lambda+\frac{1}{\tau_{life}}\right)\left(B_{1,u}+B_{1,b}+B_{3}\right)=k_{B2}B_{2},\end{split}$ (11)
+and then:
+$\begin{split}(\lambda+\frac{1}{\tau_{life}})B_{tot}&=k_{B2}B_{2}\\ (\lambda+\frac{1}{\tau_{life}}+k_{B2})B_{2}&=k_{B1}B_{1,u}\\ B_{1,b}=B_{tot}-B_{1,u}\vspace{-0.1cm}-B_{3}&=B_{tot}-B_{1,u}-\frac{k_{B2}B_{2}+k_{B4}B_{1,u}}{\lambda+k_{B3}+\frac{1}{\tau_{life}}}.\end{split}$ (12)
+From this, we recover the equivalent of the first growth law for ribosomes, in the limit $\lambda\ll k_{B2}$:
+$\begin{split}\frac{B_{1,u}}{B_{tot}}&=\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right)\left(1+\frac{\lambda}{k_{B2}}\right)B_{tot}\simeq\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right)\\ \frac{B_{2}}{B_{tot}}&=\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right)\\ \frac{B_{1,b}}{B_{tot}}&=\left[1-\frac{1}{k_{B1}}(\lambda+\frac{1}{\tau_{life}})(1+\frac{\lambda}{k_{B2}})-\frac{\lambda+\frac{1}{\tau_{life}}}{\lambda+\frac{1}{\tau_{life}}+k_{B3}}-\frac{k_{B4}(\lambda+\frac{1}{\tau_{life}})(1+\frac{\lambda}{k_{B2}})}{k_{B1}(\lambda+\frac{1}{\tau_{life}}+k_{B3})}\right],\end{split}$ (13)
+where the quantity $Q(\lambda):=\frac{1}{k_{B1}}\left(\lambda+\frac{1}{\tau_{life}}\right)\left(1+\frac{\lambda}{k_{B2}}\right)$ appears as the relative abundance of unbound ribosomes.
+We can additionally write $B_{1,b}$ in another way:
+$\begin{split}B_{1,b}&=\frac{k_{on}AB_{1,u}}{B_{tot}(\lambda+\frac{1}{\tau_{life}}+k_{off})},\end{split}$ (14)
+and compute explicitly the abundance of antibiotics:
+$A=\frac{P_{in}B_{tot}a_{ex}}{\lambda+P_{out}+\frac{k_{on}\lambda}{\lambda+\frac{1}{\tau_{life}}+k_{off}}\frac{\lambda+\frac{1}{\tau_{life}}}{k_{B1}}\left(1+\frac{\lambda}{k_{B2}}\right)}.$ (15)
+This leads to another expression for $B_{1,b}$:
+$B_{1,b}=\frac{\frac{k_{on}}{k_{B1}}P_{in}a_{ex}(\lambda+\frac{1}{\tau_{life}})B_{tot}}{\left(\lambda+\frac{1}{\tau_{life}}+k_{off}\right)\left(\lambda+P_{out}+\frac{k_{on}\lambda}{\lambda+\frac{1}{\tau_{life}}+k_{off}}\frac{\lambda+\frac{1}{\tau_{life}}}{k_{B1}}\left(1+\frac{\lambda}{k_{B2}}\right)\right)}.$ (16)
+A.1 "inhibitor-free" growth rate of the targeted network
+Without toxic agent, we obtain:
+$k_{B,1}k_{B,3}=\left(\lambda_{0}+\frac{1}{\tau_{life}}\right)\left(\lambda_{0}+\frac{1}{\tau_{life}}+k_{B,3}+k_{B,4}\right)\left(1+\frac{\lambda_{0}}{k_{B,2}}\right).$ (17)
+Here, $\lambda_{0}$ is only set by the "inhibitor-free" rates of the cell.
+As the concentration of antibiotics increases, the growth rate will be modified.
+In particular, we always have $\lambda\leq\lambda_{0}$ for bacteriostatic drugs.
+A.2 Second growth law
+To recover the second growth law, we compute:
+$\begin{split}\frac{B_{1,u}+B_{1,b}}{B_{tot}}&=1-\frac{\left(\lambda+\frac{1}{\tau_{life}}\right)\left[k_{B,1}+k_{B,4}\left(1+\frac{\lambda}{k_{B,2}}\right)\right]}{k_{B,1}\left(\lambda+\frac{1}{\tau_{life}}+k_{B,3}\right)}.\end{split}$ (18)
+In the limit of fast assembly, fast activation, and long lifetime, we find:
+$\frac{B_{1,u}+B_{1,b}}{B_{tot}}=1-\frac{\lambda}{k_{B,3}},$ (19)
+which describes the linear negative correlation found in [15].
+A.3 Self-consistent equation for the growth rate
+Let us assume $\frac{1}{\tau_{life}},k_{B4}\ll\lambda_{0},k_{B1}\ll k_{B2},k_{B3}$, which is in particular the case for ribosomes.
+This yield $\lambda_{0}\simeq k_{B1}$.
+Now equating the two equations for $B_{1,b}$:
+$\frac{P_{in}a_{ex}}{\left(\frac{k_{B1}}{k_{on}}\frac{\lambda+P_{out}}{\lambda}+\frac{\lambda}{\lambda+k_{off}}\right)}\simeq\left(1-\frac{\lambda}{\lambda_{0}}\right)\left(\lambda+k_{off}\right).$ (20)
+From this we obtain the equation defining the growth rate, also assuming fast binding $\lambda_{0}\ll k_{on}$:
+$F(\lambda)=\left(\frac{\lambda}{\lambda_{0}}\right)^{3}+\left(\frac{\lambda}{\lambda_{0}}\right)^{2}\left(\frac{k_{B,1}}{\lambda_{0}}(K_{D}+\frac{P_{out}}{k_{on}})-1\right)+\frac{\lambda}{\lambda_{0}}\left(\frac{k_{B,1}}{\lambda_{0}}K_{D}(\frac{P_{out}}{\lambda_{0}}-1)+\frac{P_{in}a_{ex}}{\lambda_{0}}-\frac{P_{out}}{\lambda_{0}}\right)-K_{D}\frac{k_{B,1}}{\lambda_{0}}\frac{P_{out}}{\lambda_{0}}=0.$ (21)
+A.3.1 Reversible limit
+In the reversible limit $P_{out},k_{off}\gg\lambda_{0}$, we obtain:
+$\frac{\lambda}{\lambda_{0}}\left(K_{D}\frac{k_{B,1}P_{out}}{\lambda_{0}^{2}}+\frac{P_{in}a_{ex}}{\lambda_{0}}\right)=K_{D}\frac{k_{B,1}P_{out}}{\lambda_{0}^{2}},$ (22)
+and therefore:
+$\lambda=\frac{\lambda_{0}}{1+\frac{\lambda_{0}P_{in}a_{ex}}{k_{B,1}K_{D}P_{out}}}.$ (23)
+We recover the result of [12] for the reversible case.
+A.3.2 Irreversible limit
+In the irreversible limit $P_{out},k_{off}\ll k_{on},\lambda_{0}$, we obtain:
+$\left(\frac{\lambda}{\lambda_{0}}\right)^{2}-\left(\frac{\lambda}{\lambda_{0}}\right)+\frac{P_{in}a_{ex}}{\lambda_{0}}=0.$ (24)
+In this case:
+$\lambda=\frac{\lambda_{0}}{2}\left(1+\sqrt{1-\frac{4P_{in}a_{ex}}{\lambda_{0}}}\right).$ (25)
+A.4 Risk induced by the toxic agent
+In the case of bacterial growth, the inhibitor is typically a bacteriostatic antibiotic.
+Antibiotics can act on the cell in various ways, for instance by binding to ribosomes [3, 2, 9, 12, 5] or by inhibiting RNA-polymerase [6].
+It seems from experiments that regardless of the mechanism of action, the effect of antibiotics on growth show similarities [11], which suggests that we could propose a general measure of the risk induced by the toxic agent.
+We propose to compare the abundance of bound active individuals $B_{1,b}$ to the abundance of unbound active individuals $B_{1,u}$ as a measure of the risk.
+Indeed, $B_{1,b}$ are either inactive or producing toxic entities (toxic proteins in the case of ribosomes [7, 30]), but represent a risk faced by the system, whereas $B_{1,u}$ are required for the cell to operate safely.
+Consequently we define a measure of the risk as:
+$\begin{split}\frac{B_{1,b}}{B_{1,u}}&=\frac{k_{B1}k_{B3}-\left(1+\frac{\lambda}{k_{B2}}\right)\left(\lambda+\frac{1}{\tau_{life}}\right)\left(\lambda+\frac{1}{\tau_{life}}+k_{B3}+k_{B4}\right)}{\left(\lambda+\frac{1}{\tau_{life}}\right)\left(1+\frac{\lambda}{k_{B2}}\right)\left(\lambda+\frac{1}{\tau_{life}}+k_{B3}\right)}\\ &\simeq\frac{k_{B1}}{\lambda}-1.\end{split}$ (26)
+The second equality corresponds to the case of ribosomes with long lifetime, fast assembly, fast activation and slow resting rates.
+We show a typical behaviour of the risk in Fig.4(a).
+In these figures, the concentration of toxic agent is rescaled by a typical concentration inspired from [12], $IC_{50}^{*}=\frac{\sqrt{K_{D}P_{out}k_{B,1}}}{P_{in}}$ and a typical growth rate $\lambda_{0}^{*}=2\sqrt{P_{out}K_{D}k_{B,1}}$.
+(a) Risk faced by the system in the presence of a toxic agent. We compare the reversible case (dotted lines) and the irreversible case (full lines).
+(b) Rescaled risk depending on the growth rate.
+We compare the reversible case (dotted lines) and the irreversible case (full lines).
+We observe a complete collapse of the curves in the reversible limit.
+The risk is rescaled by $\frac{K_{D}P_{out}}{P_{in}a_{ex}}$.
+Figure 4: Normalized risk versus antibiotic concentration
+As expected, the risk is increasing with the fraction of toxic agent.
+We also see that with this process, this measure of the risk is decreasing with $\lambda_{0}$.
+The risk increases rapidly close to $IC_{50}^{*}$, with a discontinuity at a given fraction $a_{ex,lim}$ in the irreversible case.
+This fraction can be understood as a limit concentration above which the system is significantly endangered.
+In Fig.4, we rescale the risk by $\frac{P_{in}a_{ex}}{K_{D}P_{out}}$ to obtain a collapse of the curves in the reversible limit.
+Indeed for $\frac{\lambda}{\lambda_{0}^{*}}\to 0$, the risk is equivalent to $\frac{P_{in}a_{ex}}{K_{D}P_{out}}$ in the reversible limit as can be deduced from Eq.26.
+A.5 Half-inhibition concentration
+The half-inhibition concentration $IC_{50}$ is defined as the concentration of toxic agent at which the growth rate is half its initial value $\lambda_{0}$ [12]. Therefore we have:
+$\begin{split}&IC_{50}\simeq\frac{\left(\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}+k_{off}\right)\left(\frac{\lambda_{0}}{4}+\frac{P_{out}}{2}+\frac{k_{on}\lambda_{0}}{2\lambda_{0}+\frac{4}{\tau_{life}}+4k_{off}}\frac{\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}}{k_{B1}}\left(1+\frac{\lambda_{0}}{2k_{B2}}\right)\right)}{\frac{k_{on}}{k_{B1}}P_{in}(\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}})},\end{split}$ (27)
+and in the limit of fast assembly and long lifetime:
+$\frac{IC_{50}}{IC_{50}^{*}}=\frac{1}{2}\left(\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+2K_{D}\frac{\lambda_{0}}{\lambda_{0}^{*}}\right)\left(1+\frac{\lambda_{0}}{2k_{off}}\right)+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right).$ (28)
+In the limit of long lifetime $1\gg\tau_{life}\lambda_{0}$ and fast binding $1\gg\frac{\lambda_{0}}{k_{on}}$ and fast assembly $1\gg\frac{\lambda_{0}}{k_{B,2}}$:
+$\begin{split}IC_{50}\simeq\frac{1}{4P_{in}}\left(\lambda_{0}+2k_{off}\right)\left(\frac{k_{B1}}{k_{on}}\frac{\lambda_{0}+2P_{out}}{\lambda_{0}}+\frac{\lambda_{0}}{\lambda_{0}+2k_{off}}\right).\end{split}$ (29)
+Defining $\lambda_{0}^{*}=2\sqrt{P_{out}k_{B,1}K_{D}}$, and $IC_{50}^{*}=\frac{\lambda_{0}^{*}}{2P_{in}}$ in a similar way to [12], we get:
+$\begin{split}\frac{IC_{50}}{IC_{50}^{*}}\simeq\frac{1}{2}\left(2+\frac{\lambda_{0}}{k_{off}}\right)\left(\frac{\lambda_{0}^{*}}{4}(\frac{1}{P_{out}}+\frac{2}{\lambda_{0}})+\frac{k_{off}\lambda_{0}}{\lambda_{0}^{*}(\lambda_{0}+2k_{off})}\right),\end{split}$ (30)
+for $P_{out},k_{off}\gg\lambda_{0}$:
+$\begin{split}\frac{IC_{50}}{IC_{50}^{*}}\simeq\frac{1}{2}\left(\frac{\lambda_{0}^{*}}{\lambda_{0}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\right).\end{split}$ (31)
+This is the result of [12] concerning the relationship between the "drug-free" growth rate and the half inhibition concentration.
+Appendix B List of experimental data analyzed in this work
+Chloramphenicol (Fig.5(b)) inhibits ribosome production by binding to ribosomes (preventing them from transcribing new proteins).
+Its effect on growth laws has been studied [12] as an example of bacteriostatic drug on E.Coli.
+Rifampicin (Fig.5(a)) targets RNA-polymerase by binding to RNA-polymerase [28, 29](thus inhibiting the RNA-polymerase autocatalytic cycle discussed in [16]).
+With our formalism, we also describe the effect of Triclosan (Fig.5(c)), Erythromycin (Fig.5(d)), Streptomycin (Fig.5(e)) and Kanamycin (Fig.5(f)), which have different modes of action but are all bacteriostatic drugs against E.Coli.
+Kanamycin, Streptomycin, Chloramphenicol and Erythromycin target the ribosomal autocatalytic cycle at different stages and inhibit growth [30, 3, 5, 7].
+Triclosan acts as a bacteriostatic by targeting the synthesis of fatty acids [31, 32, 33], and thus affecting the building of bacterial membranes [16].
+(a) Rifampicin targets RNA-polymerase and inhibits RNA synthesis[28]. Data from [8].
+(b) Chloramphenicol inhibits protein synthesis by binding to ribosomes [4]. Data from [8].
+(c) Triclosan targets the synthesis of fatty acids [31]. Data from [8].
+(d) Erythromycin inhibits protein synthesis by binding to ribosomal proteins [39]. Data from [8].
+(e) Streptomycin inhibits protein synthesis by binding to ribosomal proteins [12]. Data from [12].
+(f) Kanamycin inhibits protein synthesis by binding to ribosomal proteins [12]. Data from [12].
+Figure 5: Comparison with experiments for various drugs.
+In full line we show the growth rate as a function of the fraction of inhibitors.
+In dotted line we show a measure of the risk $\frac{B_{1,b}}{B_{1,u}}$.
+This measure compares the abundance of bound individuals $B_{1,b}$ to that of unbound operational individuals $B_{1,u}$ as in Eq 26.
+For ribosome-targeting drugs, this corresponds to the fraction of bound ribosomes (inhibited) to unbound ribosomes (operating).
+Unbound ribosomes are indeed required for the vital functions of the cell whereas bound ribosomes are unable to synthesize proteins.
+Appendix C General case: Arbitrary number of intermediate construction steps
+For some processes (such as the autocatalytic cycle of RNA polymerase [16]), some intermediate steps can be be significant to form mature autocatalysts $B_{1}$ as sketched on 6 (operating ribosomes or RNA-polymerase for instance).
+As an example, to form RNA-polymerase, mRNA have to be translated to resting protein subunits, that have to be activated and then assembled to form resting RNA-polymerase ($B_{N-1}$ in Fig.6, with $N=5$ in this example).
+Examples from ecology, or economy could involve slow assembly steps affecting the growth rate.
+Typically, if one sub-unit of the system is produced slowly we expect the system to be limited by this step, whereas fast assembly steps should not influence the growth rate (we can consider them to be instantaneous).
+We want to generalize the model in order to add intermediate steps.
+Below, we do this for the first cycle only, assuming $B$ is limiting as done previously.
+Figure 6: Networks with an arbitrary number of sub-steps in the first cycle.
+When adding intermediate equations for the sub-steps, the rate equations become:
+$\begin{split}\frac{dB_{1,u}}{dt}&=k_{B,N}B_{N}-k_{B,N+1}B_{1,u}-k_{on}\frac{A}{B_{tot}}B_{1,u}+k_{off}B_{1,b}-\frac{B_{1,u}}{\tau_{life}}\\ \frac{dB_{1,b}}{dt}&=k_{on}\frac{A}{B_{tot}}B_{1,u}-k_{off}B_{1,b}-\frac{B_{1,b}}{\tau_{life}}\\ \frac{dB_{2}}{dt}&=k_{B,1}B_{1,u}-k_{B,2}B_{2}\\ \vdots\\ \frac{dB_{N}}{dt}&=k_{B,N-1}B_{N-1}-k_{B,N}B_{N}+k_{B,N+1}B_{1,u}-\frac{B_{N}}{\tau_{life}}\\ \frac{dA}{dt}&=P_{in}B_{tot}a_{ex}-P_{out}A-k_{on}\frac{A}{B_{tot}}B_{1,u}+k_{off}B_{1,u}\end{split}$ (32)
+From this we obtain a system:
+$\begin{split}\left(\lambda+\frac{1}{\tau_{life}}+k_{B,N+1}+k_{on}\frac{A}{B_{tot}}\right)B_{1,u}&=k_{B,N}B_{N}+k_{off}B_{1,b}\\ \left(\lambda+\frac{1}{\tau_{life}}+k_{off}\right)B_{1,b}&=k_{on}\frac{A}{B_{tot}}B_{1,u}\\ \left(\lambda+k_{B,2}\right)B_{2}&=k_{B,1}B_{1,u}\\ \vdots\\ \left(\lambda+k_{B,N-1}\right)B_{N-1}&=k_{B,N-2}B_{N-2}\\ \left(\lambda+k_{B,N}+\frac{1}{\tau_{life}}\right)B_{N}&=k_{B,N-1}B_{N-1}+k_{B,N+1}B_{1,u}\\ \left(\lambda+P_{out}+k_{on}\frac{B_{1,u}}{B_{tot}}\right)A&=P_{in}B_{tot}a_{ex}+k_{off}B_{1,u},\end{split}$ (33)
+and if we multiply equations $3$ to $N$ together, we find:
+$B_{1,u}=\frac{\lambda+k_{B,N-1}}{k_{B,1}}\left(1+\frac{\lambda}{k_{B,2}}\right)\times\ldots\times\left(1+\frac{\lambda}{k_{B,N-2}}\right)B_{N-1}.$ (34)
+Defining $B_{tot}=B_{1,u}+B_{1,b}+B_{N}$, we obtain by summing the two first equations and the $N+1$-th:
+$B_{N-1}=\frac{1}{k_{B,N-1}}\left(\lambda+\frac{1}{\tau_{life}}\right)B_{tot},$ (35)
+and therefore, we get:
+$\frac{B_{1,u}}{B_{tot}}=\frac{1}{k_{B,1}}\left(1+\frac{\lambda}{k_{B,2}}\right)\times\ldots\times\left(1+\frac{\lambda}{k_{B,N-2}}\right)\left(1+\frac{\lambda}{k_{B,N-1}}\right)\left(\lambda+\frac{1}{\tau_{life}}\right).$ (36)
+This is the equivalent of the first growth law [13, 12, 16] in a general case.
+Indeed if all the intermediate processes are sufficiently fast $\forall n\in\{2,...,N-1\},\frac{\lambda}{k_{B,n}}\ll 1$, we recover the linear law:
+$B_{1,u}=\frac{1}{k_{B,1}}\left(\lambda+\frac{1}{\tau_{life}}\right)B_{tot}.$ (37)
+In general, $\frac{B_{1,u}}{B_{tot}}$ is a $(N-1)$-th order polynomial in $\lambda$, which we call $Q(\lambda)$.
+This polynomial is positive and increasing over $\mathbb{R}^{+}$.
+We can also express the concentration of bound individuals $B_{1,b}$:
+$\ \frac{B_{1,b}}{B_{tot}}&=\frac{k_{B,N}-Q(\lambda)(\lambda+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1})}{\lambda+\frac{1}{\tau_{life}}+k_{B,N}},\end{split}$ (38)
+we further obtain:
+$\begin{split}B_{1,u}&=Q(\lambda)B_{tot}\\ B_{1,b}&=\frac{k_{B,N}-Q(\lambda)(\lambda+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1})}{\lambda+\frac{1}{\tau_{life}}+k_{B,N}}B_{tot}\\ B_{1,b}&=\frac{k_{on}AQ(\lambda)}{\lambda+\frac{1}{\tau_{life}}+k_{off}}\\ A&=\frac{P_{in}B_{tot}a_{ex}}{\lambda+P_{out}+k_{on}Q(\lambda)\frac{\lambda+\frac{1}{\tau_{life}}}{\lambda+\frac{1}{\tau_{life}}+k_{off}}}.\end{split}$ (39)
+The second equation is obtained by writing $B_{1,b}=B_{tot}-B_{1,u}-B_{N}$.
+Equating the two equations for $B_{1,b}$, we find the general self-consistent equation on the growth rate Eq.41.
+In the absence of toxic agent, $a_{ex}=0$, the growth rate $\lambda_{0}$ is set by:
+$Q(\lambda_{0})\left(\lambda_{0}+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1}\right)=k_{B,N}.$ (40)
+As done previously, we can write a second expression for $B_{1,b}$ as proportional to the abundance of toxic agents $A$.
+Equating the two equations for $B_{1,b}$, we find a general self-consistent equation on the growth rate, which becomes equivalent to Eq.
+3 of the main text when there is only one assembly step $(N=3)$:
+$\frac{k_{on}Q(\lambda)P_{in}a_{ex}\left(\lambda+\frac{1}{\tau_{life}}+k_{B,N}\right)}{\left(\lambda+\frac{1}{\tau_{life}}+k_{off}\right)\left(\lambda+P_{out}+k_{on}Q(\lambda)\frac{\lambda+\frac{1}{\tau_{life}}}{\lambda+\frac{1}{\tau_{life}}+k_{off}}\right)}=\left(k_{B,N}-Q(\lambda)(\lambda+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1})\right).$ (41)
+In the absence of toxic agent, $a_{ex}=0$, and the growth rate $\lambda_{0}$ is set by taking the right side of the equation to be 00, which corresponds to $Q(\lambda_{0})=\frac{k_{B,N}}{\lambda_{0}+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1}}$.
+This is a generalization of the results discussed previously in the simple case.
+C.1 Reversible and irreversible limits
+C.1.1 Reversible
+In the reversible limit, $k_{off},P_{out}\gg k_{on},P_{in},...$. In this case Eq.41 becomes:
+$\begin{split}&\frac{k_{on}Q(\lambda)P_{in}a_{ex}\left(\lambda+\frac{1}{\tau_{life}}+k_{B,N}\right)}{k_{B,N}k_{off}P_{out}}=\left(1-Q(\lambda)(1+\frac{\lambda}{k_{B,N}}+\frac{1}{k_{B,N}\tau_{life}}+\frac{k_{B,N+1}}{k_{B,N}})\right),\end{split}$ (42)
+if we further assume fast assembly and long lifetime
+$\begin{split}\frac{k_{on}Q(\lambda)P_{in}a_{ex}}{k_{off}P_{out}}=\left(1-Q(\lambda)\right),\end{split}$ (43)
+and therefore:
+$\begin{split}Q(\lambda)=\frac{1}{1+\frac{K_{D}P_{in}}{P_{out}}a_{ex}}.\end{split}$ (44)
+C.1.2 Irreversible
+In the irreversible limit, $k_{off},P_{out}\ll k_{on},P_{in},...$. In this case the equation becomes:
+$\begin{split}&\frac{P_{in}a_{ex}\left(\lambda+\frac{1}{\tau_{life}}+k_{B,N}\right)}{k_{B,N}\left(\lambda+\frac{1}{\tau_{life}}\right)}=\left(1-Q(\lambda)(1+\frac{\lambda}{k_{B,N}}+\frac{1}{k_{B,N}\tau_{life}}+\frac{k_{B,N+1}}{k_{B,N}})\right),\end{split}$ (45)
+if we further assume fast assembly and long lifetime:
+$\begin{split}\frac{P_{in}a_{ex}}{\lambda}=\left(1-Q(\lambda)\right),\end{split}$ (46)
+and therefore:
+$\begin{split}\lambda Q(\lambda)-\lambda+P_{in}a_{ex}=0.\end{split}$ (47)
+C.2 Second growth law
+We can also recover a linear decreasing law between the growth rate and the ribosome fraction in the general case. With our formalism, we obtain:
+$\frac{B_{1,u}+B_{1,b}}{B_{tot}}=1-\frac{\lambda+\frac{1}{\tau_{life}}}{\lambda+\frac{1}{\tau_{life}}+k_{B,N}}-\frac{k_{B,N+1}Q(\lambda)}{\lambda+\frac{1}{\tau_{life}}+k_{B,N}}.$ (48)
+In the limit of fast assembly, fast activation, and long lifetime, we find:
+$\frac{B_{1,u}+B_{1,b}}{B_{tot}}=1-\frac{\lambda}{k_{B,N}}.$ (49)
+Again, we have a linear decreasing correlation.
+C.3 Fast assembly and long lifetime
+If we assume fast assembly and long lifetime $\forall l\in\{2,...,N\},k_{B,N+1},\frac{1}{\tau_{life}}\ll k_{B,1},\lambda_{0},\lambda\ll k_{B,l}$ we have:
+$\frac{k_{on}Q(\lambda)P_{in}a_{ex}}{\left(\lambda+k_{off}\right)\left(\lambda+P_{out}+k_{on}Q(\lambda)\frac{\lambda}{\lambda+k_{off}}\right)}=\left(1-Q(\lambda)\right),$ (50)
+and for $Q(\lambda)\simeq\frac{\lambda}{k_{B,1}}\simeq\frac{\lambda}{\lambda_{0}}$. Therefore:
+$\begin{split}F(\lambda):=&\left(\frac{\lambda}{\lambda_{0}}\right)^{3}\left(1+\frac{\lambda_{0}}{k_{on}}\right)+\left(\frac{\lambda}{\lambda_{0}}\right)^{2}\left(\frac{P_{out}}{k_{on}}+K_{D}-1-\frac{\lambda_{0}}{k_{on}}\right)+\left(\frac{\lambda}{\lambda_{0}}\right)\left(\frac{K_{D}P_{out}+P_{in}a_{ex}}{\lambda_{0}}-\frac{P_{out}}{k_{on}}-K_{D}\right)-K_{D}\frac{P_{out}}{\lambda_{0}}=0.\end{split}$ (51)
+In Fig.7, we plot the self-consistent function $F(\lambda)$, the roots of which correspond to the growth rates accessible to the system, for different values of the number of limiting steps $m$ in Fig.7.
+(a) Exact self-consistent function defining the growth rate for $m=0$.
+(b) Exact self-consistent function defining the growth rate for $m=1$.
+Figure 7: Self-consistent function, the roots of which define the growth rate. The dotted lines represent the function with increasing values of $a_{ex}$.
+Increasing the abundance of external inhibitors modifies the curvature of the self-consistent function, in particular the concave part of the function vanishes above a given concentration of toxic agents.
+For small $m$, the minimum of the function can become positive and this will induce a discontinuity in the growth rate because of the concave part of the polynomial.
+For higher values of $m$, this effect is attenuated, which smooths the behaviour of the growth rate.
+We also recover different possible behaviours for the growth rate, in particular the reversible and irreversible limits.
+As discussed in the main text, Eq.41 has two solutions in the irreversible limit, leading to two separate branches of solutions for $\lambda$.
+Taking $k_{on}\gg\lambda_{0}$, we recover the equation of [12] setting the growth rate and from which we can recover the reversible and irreversible limits.
+We can find the extrema of $F$ by studying its derivative, we find the position of the minimum of F (if it exists):
+$\lambda_{min}=\lambda_{0}\frac{\left(P_{out}-\lambda_{0}+k_{on}(K_{D}-1)\right)}{3(\lambda_{0}+k_{on})}\left(\sqrt{1-\frac{3(\lambda_{0}+k_{on})\left(k_{on}(K_{D}(\frac{P_{out}}{\lambda_{0}}-1)+P_{in}a_{ex})-P_{out}\right)}{\left(P_{out}-\lambda_{0}+k_{on}(K_{D}-1)\right)^{2}}}-1\right)$ (52)
+And the condition for irreversibility is:
+$F(\lambda_{min})\leq 0$ (53)
+C.4 Limiting intermediate steps
+Now, if we suppose that the step $n$ is considerably longer than the others,
+$\forall l\neq n,\frac{1}{\tau_{life}},k_{B,N+1}\ll k_{B,n}\ll\lambda_{0},k_{B,1}\ll k_{B,l}$
+Then:
+$\lambda_{0}^{2}=k_{B,1}k_{B,n},$ (54)
+and the growth rate of the system is $\lambda$ given by:
+$\frac{k_{on}Q(\lambda)P_{in}a_{ex}}{\left(\lambda+k_{off}\right)\left(\lambda+P_{out}+k_{on}Q(\lambda)\frac{\lambda}{\lambda+k_{off}}\right)}=\left(1-Q(\lambda)\right),$ (55)
+and $Q(\lambda)\simeq\frac{(\lambda)^{2}}{k_{B,1}k_{B,n}}=\left(\frac{\lambda}{\lambda_{0}}\right)^{2}$. Thus:
+$\frac{B_{1,b}}{B_{1,u}}=\left(\frac{\lambda_{0}}{\lambda}\right)^{2}-\left(1+\frac{\lambda}{k_{B,N}}\right),$ (56)
+and the self consistent equation becomes:
+$\begin{split}&\left(\left(\frac{\lambda}{\lambda_{0}}\right)^{2}-1\right)\left(\left(\frac{\lambda}{\lambda_{0}}\right)^{3}+\left(\lambda+P_{out}\right)\frac{\left(\lambda+k_{off}\right)}{\lambda_{0}k_{on}}\right)+\left(\frac{\lambda}{\lambda_{0}}\right)^{2}\frac{P_{in}a_{ex}}{\lambda_{0}}=0.\end{split}$ (57)
+Thus, the equation is:
+$\begin{split}F(\lambda)=\left(\frac{\lambda}{\lambda_{0}}\right)^{5}+&\left(\frac{\lambda}{\lambda_{0}}\right)^{4}\left(\frac{\lambda_{0}}{k_{on}}\right)+\left(\frac{\lambda}{\lambda_{0}}\right)^{3}\left(\frac{P_{out}}{k_{on}}+K_{D}-1\right)\\ &+\left(\frac{\lambda}{\lambda_{0}}\right)^{2}\left(\frac{K_{D}P_{out}+P_{in}a_{ex}}{\lambda_{0}}-\frac{\lambda_{0}}{k_{on}}\right)-\left(\frac{\lambda}{\lambda_{0}}\right)\left(\frac{P_{out}}{k_{on}}+K_{D}\right)-K_{D}\frac{P_{out}}{\lambda_{0}}=0.\end{split}$ (58)
+If $m$ steps are limiting in the process, we get in a similar way $Q(\lambda)=\left(\frac{\lambda}{\lambda_{0}}\right)^{m+1}$, and:
+$\begin{split}F(\lambda)=\left(\frac{\lambda}{\lambda_{0}}\right)^{2m+3}+&\left(\frac{\lambda}{\lambda_{0}}\right)^{m+3}\frac{\lambda_{0}}{k_{on}}+\left(\frac{\lambda}{\lambda_{0}}\right)^{m+2}\left(\frac{P_{out}}{k_{on}}+K_{D}-1\right)\\ &+\left(\frac{\lambda}{\lambda_{0}}\right)^{m+1}\frac{K_{D}P_{out}+P_{in}a_{ex}}{\lambda_{0}}-\left(\frac{\lambda}{\lambda_{0}}\right)^{2}\frac{\lambda_{0}}{k_{on}}-\left(\frac{\lambda}{\lambda_{0}}\right)\left(\frac{P_{out}}{k_{on}}+K_{D}\right)-K_{D}\frac{P_{out}}{\lambda_{0}}=0.\end{split}$ (59)
+The function F is shown for different cases in Fig. 7.
+C.5 Half-inhibition concentration
+By definition of the half-inhibition concentration:
+$IC_{50}=\frac{\left(\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}+k_{off}\right)\left(\frac{\lambda_{0}}{2}+P_{out}+k_{on}Q(\frac{\lambda_{0}}{2})\frac{\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}}{\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}+k_{off}}\right)\left(k_{B,N}-Q(\frac{\lambda_{0}}{2})(\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}+k_{B,N}+k_{B,N+1})\right)}{k_{on}Q(\frac{\lambda_{0}}{2})P_{in}\left(\frac{\lambda_{0}}{2}+\frac{1}{\tau_{life}}+k_{B,N}\right)},$ (60)
+in the limit of long lifetime, fast assembly and fast binding:
+$IC_{50}=\frac{\left(2^{m+1}-1\right)\left(\frac{\lambda_{0}}{2}\left(1+\frac{1}{2^{m+1}K_{D}}\right)+P_{out}\right)}{P_{in}},$ (61)
+and thus:
+$\frac{IC_{50}}{IC_{50}^{*}}=\frac{2^{m+1}-1}{2}\left(\frac{\lambda_{0}^{*}}{k_{B,1}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\left(\frac{1}{2^{m}}+K_{D}\right)\right),$ (62)
+where $\lambda_{0}^{*}=2\sqrt{\frac{\lambda_{0}^{*}}{2P_{in}}P_{out}k_{B,1}K_{D}}$ $IC_{50}^{*}=\frac{\lambda_{0}^{*}}{2P_{in}}$.
+In addition, using that $\lambda_{0}=\left(\prod_{l\text{ limiting}}k_{B,l}\right)^{\frac{1}{m}}$:
+$\frac{IC_{50}}{IC_{50}^{*}}=\frac{2^{m+1}-1}{2}\left(\frac{\lambda_{0}^{*}}{\lambda_{0}^{m+1}}\prod_{2\leq l\leq m+2\text{ limiting}}k_{B,l}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\left(\frac{1}{2^{m}}+K_{D}\right)\right).$ (63)
+For long lifetimes, fast assembly, and slow resting rate, the limit $\frac{\lambda_{0}}{\lambda_{0}^{*}}\to 0$ yields:
+$IC_{50}=(2^{m+1}-1)\left(\frac{k_{off}P_{out}}{k_{on}P_{in}}+\frac{\lambda_{0}}{2}\frac{P_{out}+k_{off}}{k_{on}P_{in}}\right),$ (64)
+and therefore:
+$\begin{split}\frac{IC_{50}}{IC_{50}^{*}}&=(2^{m+1}-1)\left(\frac{2k_{off}P_{out}}{\lambda_{0}^{*}k_{on}}+\frac{\lambda_{0}P_{out}+k_{off}}{\lambda_{0}^{*}k_{on}}\right)\\ &=(2^{m+1}-1)\left(\frac{2K_{D}P_{out}}{\lambda_{0}^{*}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}(K_{D}+\frac{P_{out}}{k_{on}})\right)\\ &=(2^{m+1}-1)\left(\sqrt{\frac{K_{D}P_{out}}{k_{B,1}}}+\frac{\lambda_{0}}{\lambda_{0}^{*}}(K_{D}+\frac{P_{out}}{k_{on}})\right).\end{split}$ (65)
+With $k_{B,1}\simeq\frac{\lambda_{0}^{m+1}}{\prod_{2\leq l\text{ limiting}}k_{B,l}}$ we have:
+$\frac{IC_{50}}{IC_{50}^{*}}=(2^{m+1}-1)\left(\frac{\lambda_{0}^{*}}{2\lambda_{0}^{m+1}}\prod_{2\leq l\text{ limiting}}k_{B,l}+\left(K_{D}+\frac{P_{out}}{k_{on}}\right)\frac{\lambda_{0}}{\lambda_{0}^{*}}\right).$ (66)
+C.6 Effect on the number of steps on the half-inhibitory concentration
+We can express this quantity in the general case from Eq.41, using its definition.
+We can also express this result for $m$ limiting steps, with $Q(\frac{\lambda_{0}}{2})\simeq\frac{1}{2^{m+1}}$, and for fast binding.
+In addition, we have $\lambda_{0}=\left(k_{B,1}\prod_{2\leq l\text{ limiting}}k_{B,l}\right)^{\frac{1}{m+1}}$, thus:
+$\frac{IC_{50}}{IC_{50}^{*}}=\frac{2^{m+1}-1}{2}\left(\frac{\lambda_{0}^{*}}{\lambda_{0}^{m+1}}\prod_{2\leq l\text{ limiting}}k_{B,l}+\frac{\lambda_{0}}{\lambda_{0}^{*}}\left(\frac{1}{2^{m}}+K_{D}\right)\right),$ (67)
+from this expression we recover the result of the simple case (or that of [12]) when $m=0$.
+We plot the rescaled half-inhibition concentration as a function of $\frac{\lambda_{0}}{\lambda_{0}^{*}}$ in Fig.
+3b of main text.
+We also notice that there is a collapse of the curves in the irreversible limit $\frac{\lambda_{0}}{\lambda_{0}^{*}}>1$.
+For long lifetimes, fast assembly, and slow resting rate, the limit $\frac{\lambda_{0}}{\lambda_{0}^{*}}\to 0$ yields:
+$\begin{split}\frac{IC_{50}}{IC_{50}^{*}}=&(2^{m+1}-1)\left(\frac{\lambda_{0}^{*}}{2\lambda_{0}^{m+1}}\prod_{2\leq l\text{ limiting}}k_{B,l}+K_{D}\left(1+\frac{(\lambda_{0}^{*})^{2}}{4k_{on}\lambda_{0}^{m+1}}\prod_{2\leq l\text{ limiting}}k_{B,l}\right)\frac{\lambda_{0}}{\lambda_{0}^{*}}\right),\end{split}$ (68)
+in the limit of fast assembly ($m=0$), this becomes $\frac{IC_{50}}{IC_{50}^{*}}=\frac{1}{2}(\frac{\lambda_{0}^{*}}{\lambda_{0}}+2K_{D}(\frac{\lambda_{0}}{\lambda_{0}^{*}}+\frac{\lambda_{0}^{*}}{4k_{on}}))$.
+We see that this expression does not depend only on the ratio $\frac{\lambda_{0}}{\lambda_{0}^{*}}$ but also on $\lambda_{0}^{*}$, which explains the slight discrepancy between the curves of Fig.
+3b (for different values of $k_{B,1}$.
+We see on Fig.
+3b of main text that it is possible to recover different regimes, with an increasing part and a decreasing part for the half-inhibition concentration in the limit of fast assembly ($m=0$).
+Adding limiting intermediate steps shifts the minimum of the parabola towards lower $\lambda_{0}$ and introduces a strong dependence on $k_{1,B}$, due to the $\lambda_{0}^{m+1}$ in Eq.68, especially for small $\lambda_{0}$ as can be seen in Eq.68.
+Noticeably, for $m=1$, the half-inhibition concentration decreases due to the limiting step for $\lambda_{0}$ small enough.
+Figure 8: $IC_{50}$ in the case $m=1$.
+Appendix D Closed compartment and long lifetime
+For a closed compartment $P_{in}=P_{out}=0$, meaning that waste only comes from the cycle itself, and if we take the long lifetime limit, the risk is:
+$\frac{B_{1,b}}{B_{1,u}}=\frac{k_{on}k_{w}Q(\lambda)}{\left(\lambda+k_{off}\right)\left(\lambda+k_{on}Q(\lambda)\frac{\lambda}{\lambda+k_{off}}\right)}.$ (69)
+In particular we can get regimes where the risk is an increasing function of the growth rate $\lambda$ as shown on Fig.9, provided $m$ is large enough.
+This regime corresponds to an accumulation of bound individuals when the growth rate is increasing, which are not diluted fast enough.
+Figure 9: Risk related to growth in a regime where risk can be increasing with $\lambda$. The full lines corresponds to a higher value of $k_{w}$ compared to the dotted lines.
+Appendix E Consequences of the growth of the first cycle on the second cycle
+For the second cycle of Fig.6 we get:
+$\begin{split}\left(\lambda+\frac{1}{\tau_{life(C)}}\right)C_{1}&=k_{C3}C_{3}-k_{C4}C_{1}\\ \left(\lambda+k_{C2}\right)C_{2}&=k_{B1}B_{1,u}\\ \left(\lambda+\frac{1}{\tau_{life(C)}}+k_{C3}\right)C_{3}&=k_{C4}C_{1}+k_{C2}C_{2}.\end{split}$ (70)
+Now if we introduce the total abundance of $C$, $C_{tot}=C_{1}+C_{3}$:
+$\begin{split}\left(\lambda+\frac{1}{\tau_{life(C)}}\right)C_{tot}&=k_{C2}C_{2}\\ \left(\lambda+k_{C2}\right)C_{2}&=k_{B1}Q(\lambda)B_{tot}\\ \left(\lambda+\frac{1}{\tau_{life(C)}}+k_{C3}\right)C_{tot}&=\left(\lambda+\frac{1}{\tau_{life(C)}}+k_{C3}+k_{C4}\right)C_{1}+k_{C2}C_{2}.\end{split}$ (71)
+We can express everything in terms of $B_{tot}$:
+$\begin{split}C_{tot}&=\frac{k_{B,1}Q(\lambda)}{\left(\lambda+\frac{1}{\tau_{life(C)}}\right)\left(1+\frac{\lambda}{k_{C2}}\right)}B_{tot},\\ C_{2}&=\frac{k_{B,1}Q(\lambda)}{\lambda+k_{C2}}B_{tot},\\ C_{1}&=\frac{k_{C3}}{\lambda+\frac{1}{\tau_{life(C)}}+k_{C3}+k_{C4}}\frac{k_{B,1}Q(\lambda)}{\left(\lambda+\frac{1}{\tau_{life(C)}}\right)\left(1+\frac{\lambda}{k_{C2}}\right)}B_{tot}.\end{split}$ (72)
+From this we see that the second cycle is affected by the toxic agent via the growth rate.
+In particular we show the effect on $C_{1}$ in Fig.10.
+We recover the distinction between the reversible and irreversible cases.
+We also observe regimes where $C_{1}$ increases with $a_{ex}$ (obtained for $\tau_{life(C)}<\tau_{life(B)}$, here the difference in lifetimes matters, because as $\lambda\to 0$, we get $\frac{C_{1}}{B_{tot}}\sim\frac{k_{C_{3}}}{\frac{1}{\tau_{life(C)}}+k_{C_{3}}+k_{C_{4}}}\frac{\tau_{life(C)}}{\tau_{life(B)}}\sim\frac{\tau_{life(C)}}{\tau_{life(B)}}$).
+In addition, we observe that for small enough values of $a_{ex}$, the relative abundance of $C_{1}$ increases with $a_{ex}$.
+In this case, the slowing down of the first cycle does not affect strongly the second cycle.
+For large concentrations of antibiotics, the first cycle is frustrated and the second one becomes limited by the need for autocatalysts of type $B$, thus leading to lower relative abundances of $C_{1}$.
+Figure 10: Fraction of autocatalysts in the second cycle when the first cycle is targeted by inhibitors.
+Appendix F Parameters for fitting the growth rates of antibiotics
+In order to recover the growth rate dependencies on drug concentration of Fig.5, we fitted our expression Eq.41 with different sets of data.
+As the number of free parameters is high, we constrained them in order to have biologically accurate values.
+From [12, 8, 40], we expect the basal growth rate $\lambda_{0}$ to be of order $1h^{-1}$.
+Similarly, the other processes of the autocatalytic cycle are expected to be of the same order and satisfying $\forall n\in\{2,...,N\},\,k_{B,N+1}<k_{B,n}<k_{B,1}$.
+In the following, we set all $k_{B,n}$ for $n\in\{2,...,N\}$ to the same value to reduce the number of free parameters.
+The binding and unbinding rates, and the influx and outflux are expected to be faster, typically ranging between $1h^{-1}$ and $1000h^{-1}$ [12, 41, 42].
+From this considerations, we allow $k_{B,1}$ to vary between $0.4h^{-1}$ and $4h^{-1}$, $P_{out}$ to vary between $0h^{-1}$ and $10^{3}h^{-1}$ and $P_{in}$ to vary between $0\mu g.mL^{-1}.h^{-1}$ and $10^{3}\mu g.mL^{-1}.h^{-1}$ to capture the effects of reversibility.
+To reduce the number of free parameters, we set $K_{D}=50\mu g/mL$ (here $K_{D}$ has a physical unit because $a_{ex}$ is a concentration).
+And $\forall n\in\{2,...,N\},k_{B,n}\in[10^{-3}h^{-1};1h^{-1}]$ and $k_{B,N+1}\in[10^{-3}h^{-1};10^{-1}h^{-1}]$.